OpenStudy (anonymous):

1. The pet store has a treat bar that sells treats for dogs. The price for 3 lb of Carob Cookies is $8.25. The price for is 5 lb of Ginger Bears is $12.75. One customer spent $21.60 on dog treats. He purchased 6 lb of Carob Cookies. How many pounds of Ginger Bears did he purchase? Write an equation for this situation.

OpenStudy (anonymous):

@Preetha plz help me

OpenStudy (anonymous):

If 3lb of Carob Cookie costs $8.25, then 6lb will cost 2*8.25=$16.5. On the other hand, if 5 lb of Ginger Bears cost $12.75, then 1lb will cost 12.25/5=$2.45. Given that the sum of the products is $21.6, we are mathematically saying that 16.5+2.45n=21.6 and n=(21.6-16.5)/2.45=2.08(pounds)

OpenStudy (anonymous):

soo... thats the answer to my problem and the equation too? Wow Thx!

OpenStudy (anonymous):

No problem.